Contribution to the construction of the CMS Tracker Endcaps (2)

O. Bouhali, B. Clerbaux, J.-P. Dewulf, J. D’Hondt, R. Goorens, J. Heyninck, S. Lowette, L. Neukermans, S. Tavernier, F. Udo, C. Vander Velde, P. Vanlaer, L. Van Lancker, C. Wastiels, J. Wickens, C. Yu

Inter-university Institute for High Energies (IIHE), Université Libre de Bruxelles and Vrije Universiteit Brussel

Belgian Physics Society, International Scientific Meeting, University of Mons-Hainaut, 25-26 May 2004 email: cms-people@hep.iihe.ac.be

4 Thermal conductivity measurement of carbon fiber frames

Responsible : Barbara

5 Deformation of the silicon sensors

The CMS detector modules consist of a sandwich of different materials glued together at room-temperature, and its operatingtemperature is -10^{o}C, so distortions will be induced because of the different expansion-coefficients. These distortions may induce several unwanted effects on the detector-modules:

• degrading of the coordinate-measurements of the module by shifting the wafers out of their original positions

• inducing of ’micro-cracks’ in the silicon-wafers at the gluing areas, leading to higher noise-ratings

• destruction of the module Figure 5: Schematic front view of the setup used for the measurement of the sensor deformation. The distortions are measured with an optical method illustrated in Figure 5. A regular array of white dots is reflected by the wafers and viewed with a ccd-camera. Any deformation of the surface translates into movement of the dots. The setup consists of the following components:

• grid: the white dots are arranged in a way that en-

sures homogeneous scanning of the investigated surface every 5mm in each direction

• ccd-camera: BW-camera with PAL-Resolution

(752x582 pixel) and a zoom-objective (12-70mm focal length)

• cooling machine: temperature range −25^{o}C to +40^{o}C, with 250W cooling power at −20^{o}C

• PC: controlling the cooling machine and reading out the camera via a frame grabber card, running under Linux

The calibration-parameters are evaluated by measuring a well known surface with the setup. In order to calibrate the setup over a wide range of distortions a special device is used consisting of a small sheet of metal which is bend using a high precision differential micrometer screw. For a silicon sensor the maximum deformation is preliminary measured to be of the order of 80μm.

6 Mounting on the petals

Responsible : Lionel

Mounting and electronic test (+ Long Term tests)

7 Outlook

Responsible : Barbara

Assembly of petals on larger structures and final integration in the CMS detector. Achievable resolution on track momenta.

References

[1] Authors, Title, Journal information.